# ImagineNav++: Prompting Vision-Language Models as Embodied Navigator through Scene Imagination

### v1

@@ III-B Future-View Imagination
To better leverage the spatial perception and reasoning capabilities of VLMs for open-vocabulary visual navigation in unknown environments, we propose a future-view imagination model, which is composed of a Where2Imagine module followed by a novel view synthesis (NVS) model. As illustrated in Figure 2, the Where2Imagine predicts the relative pose ($\Delta x,\Delta y,\Delta\theta$) of a potential next waypoint from the current RGB observation, where $\Delta x$, $\Delta y$ and $\Delta\theta$ denote lateral displacement, longitudinal displacement, and viewing angle change, respectively. This predicted pose is subsequently fed into the NVS module to generate a corresponding imagined observation. Recent years have witnessed remarkable progress in NVS, with advanced methods ranging from few-shot 3D rendering techniques [sargent2024zeronvs, wimbauerbehind, cao2023scenerf] to generative models, particularly diffusion models [yu2024polyoculussimultaneousmultiviewimagebased, tseng2023consistentviewsynthesisposeguided, yu2023longtermphotometricconsistentnovel], which have demonstrated powerful image synthesis capabilities across diverse domains. In this work, we adopt a pre-trained diffusion model “Polyoculus” [yu2024polyoculussimultaneousmultiviewimagebased] for future-view imagination owing to its capacity to generate perceptually consistent and high-fidelity novel views from a single RGB image and a relative pose.
@@
Specifically, to equip the Where2Imagine module with human-like spatial intuition for waypoint prediction, we train a ResNet-18 model [he2016resent] from scratch using human demonstration data sourced from the Habitat-Web project [ramrakhya2022habitat]. The data comprises 80k ObjectNav and 12k Pick&Place trajectories collected via a virtual teleoperation system on Amazon Mechanical Turk, capturing natural human navigation behaviors in indoor environments. A key insight derived from these demonstrations is that humans consistently favor directions toward semantically meaningful structures (e.g., doors) that facilitate exploration, as illustrated in Figure 3. By learning from these demonstrations, the model acquires semantically-informed navigational preferences, enabling it to predict waypoints that are both geometrically plausible and semantically meaningful, thereby supporting efficient and contextually-aware navigation. We reformulated the human demonstration trajectories into a paired dataset $\{(O_{t},P_{t+T})\}$, where $P_{t+T}=(\Delta x,\Delta y,\Delta\theta)$ denotes the ground-truth relative waypoint pose with respect to the agent’s current frame $O_{t}$. Here, $T$ defines a predictive temporal horizon that captures how humans anticipate and navigate through semantically meaningful spatial scales. To enhance learning efficacy, we apply a depth-based filter to exclude observations with limited semantic content (e.g., close-up views of blank walls):
@@ III-C Selective Foveation Memory
In complex indoor navigation scenarios, keyframes serve as compact yet informative representations that capture salient spatial and contextual information across temporal sequences. They reduce data redundancy and preserve essential perceptual cues, thereby enhancing memory efficiency while supporting long-term temporal reasoning. However, reliably extracting such keyframes from observational data remains challenging, particularly in identifying semantically meaningful segments within extended visual streams. Traditional keyframe extraction methods often rely on hand-crafted features such as color histograms [VSUMM2011, Ejaz2012, Wu2017] for clustering, which struggles to capture deep semantic similarities between frames. In contrast, modern deep learning approaches can model complex spatio-temporal dependencies but are typically heavily supervised, requiring extensive manual annotations [Zhao2018, Rochan2018, Elfeki2019, Jiang2022]. This motivates our exploration of DINOv2 [DINOv2] for keyframe extraction in a zero-shot manner, where selection is driven primarily by inter-frame visul similarity. DINOv2’s self-supervised pre-training on a large-scale and diverse image corpus enables it to learn unified visual representations that inherently encode holistic scene structure. Consequently, the model exhibits a pronounced sensitivity to scene-level variations, making it particularly suitable for identifying semantically representative keyframes without task-specific training.
@@ III-D High-level Planning
The high-level planning module leverages the spatial awareness and commonsense reasoning capabilities of the VLM to select the direction most conducive to locating the navigation target. We employ GPT-4o-mini [openai2024gpt4omini] as the high-level planner due to its favorable balance between reasoning performance and practical efficiency. To assist GPT-4o-mini in decision making, we designed a simple prompt template, requiring the VLM to summarize its choice in a JSON format containing { ‘Reason’, ‘Choice’ }. This format allows for a clear understanding of the VLM’s reasoning process. As illustrated in Figure 2, the VLM is provided with synthesized observations of potential future waypoints, historical memory, and the navigation goal. Guided by the hierarchical prompt, it analyzes the semantic content of each view to select the optimal exploration direction, returning its decision in a structured format. The complete prompt can be found in section IV. This integration of imagined future observations and historical keyframes as as visual prompts equips our ImagineNav++ with significantly enhanced spatial reasoning and long-term decision-making capabilities. First, VLMs are more skilled at handling multiple-choice decision tasks compared to 3D geometry question answering (i.e., directly inferring the 3D coordinates of next waypoints). Furthermore, the introduction of the imagined future observations augment the VLM’s scene understanding by providing rich contextual information about distant or visually unclear objects, while the historical keyframes maintain a coherent representation of the environment’s state over time. This complementary mechanism effectively mitigates perceptual uncertainty and facilitates robust long-horizon reasoning. The proposed imagination-guided reasoning procedure–integrating waypoint imagination and high-level planning–operates on a periodic cycle: it is triggered once every $T$ steps, proposing a new sub-goal only after the low-level controller has successfully completed navigation to the current target. This design ensures an effective balance between computational efficiency and robust long-horizon navigation performance.
@@ III-E Low-level Controller
Following waypoint selection by the high-level planner, the low-level controller executes a Point Goal Navigation (PointNav) strategy to reach each designated target. In contrast to ObjectNav, which depends on semantic cues from the environment, PointNav operates purely on spatial perception, using relative displacement commands ($\Delta x$, $\Delta y$) without requiring semantic understanding. Multiple established methods exist for PointNav implementation [yang2023iplanner, roth2024viplanner, wijmans2022ver, liang2024mtg]. In our framework, we adopt Variable Experience Rollout (VER [wijmans2022ver]) as the underlying policy for action selection at each navigation step. VER combines the advantages of synchronous and asynchronous reinforcement learning paradigms, leading to improved training efficiency and sample utilization. As a result, the agent exhibits enhanced adaptability and generalization performance in novel and complex environments.
TABLE I: ImagineNav++: Comparison with previous work on object-goal navigation. The Where2Imagine model with TT=11, utilizing ResNet-18 trained from scratch and GPT-4o-mini as the VLM, was evaluated over 200 epochs on the Gibson, HM3D, and HSSD datasets. ImagineNav++ uses NVS model to generate novel view images, while ImagineNav++(Oracle) uses real images of the candidate points to facilitate spatial reasoning.
@@
FBE [topiwala2018frontier] 64.3 28.3 33.7 15.3 36.0 17.7
SemExp [chaplot2020object] 71.7 39.6 37.9 18.8 - -
Habitat-Web [ramrakhya2022habitat] - - 41.5 16.0 - -
OVRL [yadav2023offline] - - 62.0 26.8 - -
ESC [Zhou2023ESCEW] - - 39.2 22.3 - -
VoroNav [wu2024voronav] - - 42.0 26.0 41.0 23.2
VLFM [yokoyama2024vlfm] 84.0 52.2 52.5 30.4 - -
Goat [chang2024goatthing] - - 50.6 24.1 - -
SG-Nav [yin2024sgnavonline3dscene] - - 54.0 24.9 - -
UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] - - 54.5 25.1 - -
ZSON [majumdar2022zson] 31.3 12.0 25.5 12.6 - -
PixNav [cai2023bridgingzeroshotobjectnavigation] - - 37.9 20.5 - -
PSL [sun2024prioritizedsemanticlearningzeroshot] - - 42.4 19.2 - -
ImagineNav [zhao2025imaginenavpromptingvisionlanguagemodels] - - 53.0 23.8 51.0 24.9
@@ IV-A Experimental Setup and Evaluation metrics
We evaluate the effectiveness and navigation efficiency of our proposed ImagineNav++ using the Habitat v3.0 simulator [puig2023habitat] on both Object-Goal Navigation (ObjectNav) and Instance-Image-Goal Navigation (InsINav). For ObjectNav, we conduct experiments on three widely adopted datasets: Gibson [xia2018gibson], HM3D [ramakrishnan2021hm3dc] and HSSD [khanna2023habitatsyntheticscenesdataset]. Specifically, We employ the ObjectNav training and validation splits for Gibson, as introduced in SemExp [chaplot2020object], covering 27 scenes and 6 object categories. The HM3D dataset, employed in the Habitat 2022 ObjectNav Challenge, contains 2,000 validation episodes across 20 unique environments and 6 object categories. The recently introduced HSSD dataset comprises 40 high-quality synthetic scenes with 1,200 validation episodes covering 6 object categories. For InsINav, we follow established protocols from prior works [sun2024prioritizedsemanticlearningzeroshot, yin2025unigoaluniversalzeroshotgoaloriented] to evaluate performance on the HM3D dataset [ramakrishnan2021hm3dc]. The experimental configuration adheres to the ObjectNav Challenge 2023 guidelines [habitatchallenge2023]. To support the data collection for the Where2Imagine module, we utilize human demonstration trajectories from the MP3D dataset [Matterport3D] within the habitat-web project, using a camera height of 0.88 m and a horizontal field of view (HFOV) of 79°.
We evaluate visual navigation performance using two standard metrics: Success Rate (SR) and Success weighted by Path Length (SPL) [anderson2018evaluation]. SR is defined as the proportion of episodes in which the agent executes the STOP action within 1 meter of the target object. SPL is computed as:
@@ IV-B Baselines
For ObjectNav, we conducted a comparative analysis of non-zero-shot and zero-shot object navigation methods to rigorously evaluate the superiority of our ImagineNav++ framework. FBE [topiwala2018frontier] pioneered a frontier-based exploration strategy, emphasizing the boundaries between explored and unexplored regions. SemExp [chaplot2020object] advanced this concept by implementing goal-directed semantic exploration through the construction of semantic maps. In addition, we examined non-mapping closed-set visual navigation baselines, including approaches based on imitation learning [ramrakhya2022habitat] and visual representation learning [yadav2023offline]. For zero-shot object navigation, we evaluate several mapping-based baselines  [Zhou2023ESCEW, wu2024voronav, yokoyama2024vlfm, yin2024sgnavonline3dscene, yin2025unigoaluniversalzeroshotgoaloriented, chang2024goatthing]. Among these, [Zhou2023ESCEW, wu2024voronav, yokoyama2024vlfm, chang2024goatthing] maintains a semantic 2D map of the scene and leverages semantic knowledge to facilitate navigation towards target objects, while SG-Nav [yin2024sgnavonline3dscene] and UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] represent the observed scene with 3D scene graph and leverage LLMs for explicit graph-based reasoning. Additionally, we also investigate RGB-based non-mapping navigation baselines, including ZSON [majumdar2022zson], PSL[sun2024prioritizedsemanticlearningzeroshot] and PixNav [cai2023bridgingzeroshotobjectnavigation]. Specifically, ZSON and PSL employ the CLIP model [radford2021learning] to embed both target images and object goals into a unified semantic space, enabling the training of a semantic goal-driven navigation agent; PixNav, on the other hand, utilizes pixel-level guidance provided by VLMs and LLMs to achieve pixel-accurate navigation.
@@
For InsINav, we compare with the supervised methods Krantz et al. [krantz2022instance], OVRL-v2 [yadav2023ovrl], and zero-shot methods Mod-IIN [krantz2023navigatingobjectsspecifiedimages], Goat [chang2024goatthing], UniGoal [yin2025unigoaluniversalzeroshotgoaloriented], and PSL [sun2024prioritizedsemanticlearningzeroshot]. Krantz et al. [krantz2022instance] pioneered the InsINav task and established an end-to-end RL baseline based on proximal policy optimization (PPO [schulman2017proximal]) and variable experience rollout (VER [wijmans2022ver]). OVRL-v2 [yadav2023ovrl] employs an architecture that integrates ViT with LSTM, achieving end-to-end navigation learning through self-supervised visual pre-training, while OVRL-v2-IIN is specifically fine-tuned for InsINav task using the protocol from OVRL-v2 [yadav2023ovrl]. Among these zero-shot methods, Mod-IIN [krantz2023navigatingobjectsspecifiedimages] is specifically tailored for InsINav task, which re-identifies the goal instance in egocentric vision using feature-matching and localize the goal instance by projecting matched features to a map. Goat [chang2024goatthing] enables lifelong navigation by constructing an instance-aware semantic memory, utilizing CLIP [radford2021learning] and SuperGlue [sarlin2020superglue] for multimodal goal matching, and integrating frontier exploration with classical path planning. In contrast, UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] achieves universal navigation by aligning diverse goals and scene graphs into a shared space through a unified graph representation and leveraging graph matching to dynamically guide its exploration.
@@ IV-C Numerical Evaluations on ObjectNav Task
Table I presents a comparative analysis of the proposed ImagineNav++ framework against prior state-of-the-art methods. On the Gibson dataset, our ImagineNav++ achieves state-of-the-art performance under the open-vocabulary and mapless setting. While VLFM [yokoyama2024vlfm] reports a higher SR (84.0%) on Gibson, its reliance on real-time dense semantic-geometric mapping makes it vulnerable to cumulative localization drift, particularly in large-scale environments, which limits practical robustness and scalability. In contrast, ImagineNav++ attains a competitive 72.4% SR without explicit mapping, demonstrating greater system robustness and deployment flexibility while retaining open-vocabulary understanding. This advantage is convincingly validated on the larger and more complex HM3D and HSSD dataset. Specifically, on HM3D, our ImagineNav++ maintains leading performance among open-vocabulary methods, achieving a 58.5% SR and a 26.6% SPL, surpassing map-based VLFM [yokoyama2024vlfm] and graph-based Unigoal [yin2025unigoaluniversalzeroshotgoaloriented] by 6.0% and 4.0% respectively in success rate. Moreover, ImagineNav++ achieves the highest SR and SPL on the HSSD dataset. The above observations indicate that our ImagineNav++ demonstrates outstanding navigation performance across various settings, while maintaining low storage and computational complexities. Furthermore, since the pretrained NVS is directly employed without finetunned on the Gibson, HM3D and HSSD datasets, we see a disparity between the quality of images generated by the NVS model and real images, limiting the capability of our model to some extent. To explore the upper limits of our framework, we instead use real panoramic images-specifically, the observation at the pose predicted by the Where2Imagine module-as visual prompts for the VLM model. Notably, both the success rate and SPL exhibit obvious improvements, obtaining 77.1%, 62.5% and 67.5% at SR respectively on Gibson, H3MD and HSSD benchmarks, which further demonstrates the superiority of our imagination-based navigation framework.
TABLE II: ImagineNav++: Comparison with previous work on instance-image-goal navigation. The Where2Imagine model with TT=11, utilizing ResNet-18 trained from scratch and GPT-4o-mini as the VLM, was evaluated over 200 epochs on the HM3D dataset.
@@
Krantz et al. [krantz2022instance] 8.3 3.5
OVRL-v2-IIN [yadav2023ovrl] 24.8 11.8
Mod-IIN [krantz2023navigatingobjectsspecifiedimages] 56.1 23.3
Goat [chang2024goatthing] 50.6 24.1
UniGoal [yin2025unigoaluniversalzeroshotgoaloriented] 60.2 23.7
PSL [sun2024prioritizedsemanticlearningzeroshot] 23.0 11.4

### v2

@@ III-B Future-View Imagination
To better leverage the spatial perception and reasoning capabilities of VLMs for open-vocabulary visual navigation in unknown environments, we propose a future-view imagination model, which is composed of a Where2Imagine module followed by a novel view synthesis (NVS) model. As illustrated in Figure 2, the Where2Imagine predicts the relative pose ($\Delta x,\Delta y,\Delta\theta$) of a potential next waypoint from the current RGB observation, where $\Delta x$, $\Delta y$ and $\Delta\theta$ denote lateral displacement, longitudinal displacement, and viewing angle change, respectively. This predicted pose is subsequently fed into the NVS module to generate a corresponding imagined observation. Recent years have witnessed remarkable progress in NVS, with advanced methods ranging from few-shot 3D rendering techniques [60, 73, 7] to generative models, particularly diffusion models [87, 70, 88], which have demonstrated powerful image synthesis capabilities across diverse domains. In this work, we adopt a pre-trained diffusion model “Polyoculus” [87] for future-view imagination owing to its capacity to generate perceptually consistent and high-fidelity novel views from a single RGB image and a relative pose.
@@
Specifically, to equip the Where2Imagine module with human-like spatial intuition for waypoint prediction, we train a ResNet-18 model [25] from scratch using human demonstration data sourced from the Habitat-Web project [57]. The data comprises 80k ObjectNav and 12k Pick&Place trajectories collected via a virtual teleoperation system on Amazon Mechanical Turk, capturing natural human navigation behaviors in indoor environments. A key insight derived from these demonstrations is that humans consistently favor directions toward semantically meaningful structures (e.g., doors) that facilitate exploration, as illustrated in Figure 3. By learning from these demonstrations, the model acquires semantically-informed navigational preferences, enabling it to predict waypoints that are both geometrically plausible and semantically meaningful, thereby supporting efficient and contextually-aware navigation. We reformulated the human demonstration trajectories into a paired dataset $\{(O_{t},P_{t+T})\}$, where $P_{t+T}=(\Delta x,\Delta y,\Delta\theta)$ denotes the ground-truth relative waypoint pose with respect to the agent’s current frame $O_{t}$. Here, $T$ defines a predictive temporal horizon that captures how humans anticipate and navigate through semantically meaningful spatial scales. To enhance learning efficacy, we apply a depth-based filter to exclude observations with limited semantic content (e.g., close-up views of blank walls):
@@ III-C Selective Foveation Memory
In complex indoor navigation scenarios, keyframes serve as compact yet informative representations that capture salient spatial and contextual information across temporal sequences. They reduce data redundancy and preserve essential perceptual cues, thereby enhancing memory efficiency while supporting long-term temporal reasoning. However, reliably extracting such keyframes from observational data remains challenging, particularly in identifying semantically meaningful segments within extended visual streams. Traditional keyframe extraction methods often rely on hand-crafted features such as color histograms [4, 15, 74] for clustering, which struggles to capture deep semantic similarities between frames. In contrast, modern deep learning approaches can model complex spatio-temporal dependencies but are typically heavily supervised, requiring extensive manual annotations [97, 58, 16, 30]. This motivates our exploration of DINOv2 [51] for keyframe extraction in a zero-shot manner, where selection is driven primarily by inter-frame visual similarity. DINOv2’s self-supervised pre-training on a large-scale and diverse image corpus enables it to learn unified visual representations that inherently encode holistic scene structure. Consequently, the model exhibits a pronounced sensitivity to scene-level variations, making it particularly suitable for identifying semantically representative keyframes without task-specific training.
@@ III-D High-level Planning
The high-level planning module leverages the spatial awareness and commonsense reasoning capabilities of the VLM to select the direction most conducive to locating the navigation target. We employ GPT-4o-mini [50] as the high-level planner due to its favorable balance between reasoning performance and practical efficiency. To assist GPT-4o-mini in decision making, we designed a simple prompt template, requiring the VLM to summarize its choice in a JSON format containing { ‘Reason’, ‘Choice’ }. This format allows for a clear understanding of the VLM’s reasoning process. As illustrated in Figure 2, the VLM is provided with synthesized observations of potential future waypoints, historical memory, and the navigation goal. Guided by the hierarchical prompt, it analyzes the semantic content of each view to select the optimal exploration direction, returning its decision in a structured format. The complete prompt can be found in section IV. This integration of imagined future observations and historical keyframes as visual prompts equips our ImagineNav++ with significantly enhanced spatial reasoning and long-term decision-making capabilities. First, VLMs are more skilled at handling multiple-choice decision tasks compared to 3D geometry question answering (i.e., directly inferring the 3D coordinates of next waypoints). Furthermore, the introduction of the imagined future observations augment the VLM’s scene understanding by providing rich contextual information about distant or visually unclear objects, while the historical keyframes maintain a coherent representation of the environment’s state over time. This complementary mechanism effectively mitigates perceptual uncertainty and facilitates robust long-horizon reasoning. The proposed imagination-guided reasoning procedure–integrating waypoint imagination and high-level planning–operates on a periodic cycle: it is triggered once every $T$ steps, proposing a new sub-goal only after the low-level controller has successfully completed navigation to the current target. This design ensures an effective balance between computational efficiency and robust long-horizon navigation performance.
@@ III-E Low-level Controller
Following waypoint selection by the high-level planner, the low-level controller executes a Point Goal Navigation (PointNav) strategy to reach each designated target. In contrast to ObjectNav, which depends on semantic cues from the environment, PointNav operates purely on spatial perception, using relative displacement commands ($\Delta x$, $\Delta y$) without requiring semantic understanding. Multiple established methods exist for PointNav implementation [82, 59, 72, 40]. In our framework, we adopt Variable Experience Rollout (VER [72]) as the underlying policy for action selection at each navigation step. VER combines the advantages of synchronous and asynchronous reinforcement learning paradigms, leading to improved training efficiency and sample utilization. As a result, the agent exhibits enhanced adaptability and generalization performance in novel and complex environments.
TABLE I: ImagineNav++: Comparison with previous work on object-goal navigation. The Where2Imagine module with TT=11, utilizing ResNet-18 trained from scratch and GPT-4o-mini as the VLM, was evaluated over 200 epochs on the Gibson, HM3D, and HSSD datasets. ImagineNav++ uses NVS model to generate novel view images, while ImagineNav++(Oracle) uses real images of the candidate points to facilitate spatial reasoning.
@@
FBE [67] 64.3 28.3 33.7 15.3 36.0 17.7
SemExp [10] 71.7 39.6 37.9 18.8 - -
Habitat-Web [57] - - 41.5 16.0 - -
OVRL [80] - - 62.0 26.8 - -
ESC [100] - - 39.2 22.3 - -
VoroNav [76] - - 42.0 26.0 41.0 23.2
VLFM [85] 84.0 52.2 52.5 30.4 - -
Goat [9] - - 50.6 24.1 - -
SG-Nav [83] - - 54.0 24.9 - -
UniGoal [84] - - 54.5 25.1 - -
ZSON [47] 31.3 12.0 25.5 12.6 - -
PixNav [6] - - 37.9 20.5 - -
PSL [64] - - 42.4 19.2 - -
ImagineNav [98] - - 53.0 23.8 51.0 24.9
@@ IV-A Experimental Setup and Evaluation Metrics
We evaluate the effectiveness and navigation efficiency of our proposed ImagineNav++ using the Habitat v3.0 simulator [52] on both Object-Goal Navigation (ObjectNav) and Instance-Image-Goal Navigation (InsINav). For ObjectNav, we conduct experiments on three widely adopted datasets: Gibson [77], HM3D [56] and HSSD [32]. Specifically, we employ the ObjectNav training and validation splits for Gibson, as introduced in SemExp [10], covering 27 scenes and 6 object categories. The HM3D dataset, employed in the Habitat 2022 ObjectNav Challenge, contains 2,000 validation episodes across 20 unique environments and 6 object categories. The recently introduced HSSD dataset comprises 40 high-quality synthetic scenes with 1,200 validation episodes covering 6 object categories. For InsINav, we follow established protocols from prior works [64, 84] to evaluate performance on the HM3D dataset [56]. The experimental configuration adheres to the ObjectNav Challenge 2023 guidelines [78]. To support the data collection for the Where2Imagine module, we utilize human demonstration trajectories from the MP3D dataset [8] within the habitat-web project, using a camera height of 0.88 m and a horizontal field of view (HFOV) of 79°.
We evaluate visual navigation performance using two standard metrics: Success Rate (SR) and Success weighted by Path Length (SPL) [3]. SR is defined as the proportion of episodes in which the agent executes the STOP action within 1 meter of the target object. SPL is computed as:
@@ IV-B Baselines
For ObjectNav, we conducted a comparative analysis of non-zero-shot and zero-shot object navigation methods to rigorously evaluate the superiority of our ImagineNav++ framework. FBE [67] pioneered a frontier-based exploration strategy, emphasizing the boundaries between explored and unexplored regions. SemExp [10] advanced this concept by implementing goal-directed semantic exploration through the construction of semantic maps. In addition, we examined non-mapping closed-set visual navigation baselines, including approaches based on imitation learning [57] and visual representation learning [80]. For zero-shot object navigation, we evaluate several mapping-based baselines  [100, 76, 85, 83, 84, 9]. Among these, [100, 76, 85, 9] maintains a semantic 2D map of the scene and leverages semantic knowledge to facilitate navigation towards target objects, while SG-Nav [83] and UniGoal [84] represent the observed scene with 3D scene graph and leverage LLMs for explicit graph-based reasoning. Additionally, we also investigate RGB-based non-mapping navigation baselines, including ZSON [47], PSL[64] and PixNav [6]. Specifically, ZSON and PSL employ the CLIP model [53] to embed both target images and object goals into a unified semantic space, enabling the training of a semantic goal-driven navigation agent; PixNav, on the other hand, utilizes pixel-level guidance provided by VLMs and LLMs to achieve pixel-accurate navigation.
@@
For InsINav, we compare with the supervised methods Krantz et al. [35], OVRL-v2 [79], and zero-shot methods Mod-IIN [34], Goat [9], UniGoal [84], and PSL [64]. Krantz et al. [35] pioneered the InsINav task and established an end-to-end RL baseline based on proximal policy optimization (PPO [62]) and variable experience rollout (VER [72]). OVRL-v2 [79] employs an architecture that integrates ViT with LSTM, achieving end-to-end navigation learning through self-supervised visual pre-training, while OVRL-v2-IIN is specifically fine-tuned for InsINav task using the protocol from OVRL-v2 [79]. Among these zero-shot methods, Mod-IIN [34] is specifically tailored for InsINav task, which re-identifies the goal instance in egocentric vision using feature-matching and localizes the goal instance by projecting matched features to a map. Goat [9] enables lifelong navigation by constructing an instance-aware semantic memory, utilizing CLIP [53] and SuperGlue [61] for multimodal goal matching, and integrating frontier exploration with classical path planning. In contrast, UniGoal [84] achieves universal navigation by aligning diverse goals and scene graphs into a shared space through a unified graph representation and leveraging graph matching to dynamically guide its exploration.
@@ IV-C Numerical Evaluations on ObjectNav Task
Table I presents a comparative analysis of the proposed ImagineNav++ framework against prior state-of-the-art methods. On the Gibson dataset, our ImagineNav++ achieves state-of-the-art performance under the open-vocabulary and mapless setting. While VLFM [85] reports a higher SR (84.0%) on Gibson, its reliance on real-time dense semantic-geometric mapping makes it vulnerable to cumulative localization drift, particularly in large-scale environments, which limits practical robustness and scalability. In contrast, ImagineNav++ attains a competitive 72.4% SR without explicit mapping, demonstrating greater system robustness and deployment flexibility while retaining open-vocabulary understanding. This advantage is convincingly validated on the larger and more complex HM3D and HSSD dataset. Specifically, on HM3D, our ImagineNav++ maintains leading performance among open-vocabulary methods, achieving a 58.5% SR and a 26.6% SPL, surpassing map-based VLFM [85] and graph-based Unigoal [84] by 6.0% and 4.0% respectively in success rate. Moreover, ImagineNav++ achieves the highest SR and SPL on the HSSD dataset. The above observations indicate that our ImagineNav++ demonstrates outstanding navigation performance across various settings, while maintaining low storage and computational complexities. Furthermore, since the pretrained NVS is directly employed without finetunned on the Gibson, HM3D and HSSD datasets, we see a disparity between the quality of images generated by the NVS model and real images, limiting the capability of our model to some extent. To explore the upper limits of our framework, we instead use real panoramic images-specifically, the observation at the pose predicted by the Where2Imagine module-as visual prompts for the VLM model. Notably, both the success rate and SPL exhibit obvious improvements, obtaining 77.1%, 62.5% and 67.5% at SR respectively on Gibson, H3MD and HSSD benchmarks, which further demonstrates the superiority of our imagination-based navigation framework.
TABLE II: ImagineNav++: Comparison with previous work on instance-image-goal navigation. The Where2Imagine module with TT=11, utilizing ResNet-18 trained from scratch and GPT-4o-mini as the VLM, was evaluated over 200 epochs on the HM3D dataset.
@@
Krantz et al. [35] 8.3 3.5
OVRL-v2-IIN [79] 24.8 11.8
Mod-IIN [34] 56.1 23.3
Goat [9] 50.6 24.1
UniGoal [84] 60.2 23.7
PSL [64] 23.0 11.4